OpenStudy (anonymous):

Verify this trigonometric equation...- tan^2x + sec^2x = 1?

OpenStudy (anonymous):

- tan^2x + sec^2x=1 or sec^2x- tan^2x =1 consider the following triangle ABC in which angle B=90 degrees and angle C=x

OpenStudy (anonymous):

1 + tan^2u = sec^2u is an identity so is that part of it?

OpenStudy (anonymous):

@dpasingh ?

OpenStudy (anonymous):

sorry some problem had arisen with my pc.

OpenStudy (anonymous):



OpenStudy (anonymous):

I am a little confused.. how can a triangle be used to verify?

OpenStudy (anonymous):

sec x= AC/BC and tan x = AB/BC i.e. \[\sec ^{2}x = [\frac{AC}{BC}] ^2 \] \[\tan ^{2}x = [\frac{AB}{BC}] ^2 \] \[\sec^2x-\tan^2x= \frac{AC^2}{BC^2}-\frac{AB^2}{BC^2}=\frac{AC^2-AB^2}{BC^2}=\frac{BC^2}{BC^2}=1\] Hence \[\sec^2x-\tan^2x= 1\] Thus proved Since by Pythagoras theorem: \[AC^2= AB^2+BC^2 \rightarrow AC^2-AB^2= BC^2\rightarrow BC^2=AC^2-AB^2\]

OpenStudy (raden):

yes, 1 + tan^2x = sec^2 x - tan^2x + 1 + tan^2x = 1

OpenStudy (anonymous):

Thank You guys! Really helpful!

OpenStudy (raden):

yw